OpenStudy (anonymous):

solve the following equation for the variable x 3/x-2 + 1/x+2 = -6x/x^2-4

OpenStudy (anonymous):

first, add the two fractions together by finding a common denominator. so it should look like 3(x+2)+(x-2)/(x-2)(x+2). notice that (x-2)(x+2)=x^2-4. so the left hand side looks like 3(x+2)+(x-2)/(x^2-4). the right hand side denominator should cancel with the left hand denominator. then its just a matter of solving 3(x+2)+(x-2)=-6x